Frank is standing on frictionless ice when Dave tosses a 

4.7 kg pumpkin to him. Frank catches the pumpkin and 

slides off in the same direction as the pumpkin with a 

combined speed of 0.16 m/s. If Frank’s mass is 61.0 kg 

what was the velocity of the pumpkin?


1
Expert's answer
2021-03-26T20:10:55-0400

Let the direction to Frank be the positive direction. Then, we can write the law of conservation of momentum:


"m_pv_p+m_Fv_F=(m_p+m_F)v_c,""m_pv_p+0=(m_p+m_F)v_c,""v_p=\dfrac{(m_p+m_F)v_c}{m_p},""v_p=\dfrac{(4.7\ kg+61\ kg)\cdot0.16\ \dfrac{m}{s}}{4.7\ kg}=2.24\ \dfrac{m}{s}."